the measures of two complementary angles are 16z - 9 and 4z + 3. find the measures of the angles
Answer by rfer(16322)

16z-9+4z+3=90
20z=96
z=4.8 degrees
16z-9=67.8 degrees
4z+3=22.2 degrees